OpenStudy (anonymous):

Does anyone know how to do the "P.I.A.N.3"? Performing four of the arithmetic operations using like and unlike radical terms and expressing the result in the simplest form? Please help :'l

OpenStudy (anonymous):

\[3\sqrt{5}+5\sqrt{5}-3\sqrt{5}\]

OpenStudy (anonymous):

Answer choices are.. 11 radical 5, 5 radical 5, radical 25, & 25

OpenStudy (anonymous):

Treat the \[\sqrt{5}\] as you would an x or x^2 ... in a polynomial

OpenStudy (anonymous):

So.. would the answer be 5 radical 5?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thanks I think I get it now.

OpenStudy (anonymous):

Great : )

OpenStudy (anonymous):

Wait how about if it's \[-8\sqrt{7}-\sqrt{36}-5\sqrt{63}\]

OpenStudy (anonymous):

Would it be \[-3\sqrt{92}?\]

OpenStudy (jennychan12):

-8√7 - 6 -15√7

OpenStudy (jennychan12):

-8√7 - 6 -15√7 √7 (-8-15) -6 i got -23√7 -6

OpenStudy (anonymous):

I don't get it :l

OpenStudy (jennychan12):

√63 simplifies to 3√7

OpenStudy (anonymous):

Okay thanks, how about this one. \[\frac{ 4\sqrt{3}-\sqrt{3} }{ 6 }\]

OpenStudy (anonymous):

Would it be 24?

OpenStudy (jennychan12):

√3( 4-1) / 6 sorry it's √3/2

OpenStudy (anonymous):

My answer choices are radical 3 over 2, 2 over 3, 2, 6, and 0. So the answer would be the first choice right?

OpenStudy (jennychan12):

yep

OpenStudy (anonymous):

How about Darin simplified \[5\sqrt{5}+2\sqrt{5}\] and got 15.7 Martha simplified the same expression and got 50.Use a calculator to determine who got the correct answer?